OpenStudy (anonymous):

POint X is located at (2,4), and y is located at (-6,10). FInd the y value of the point that lies halfway between X and Y. 2 7 -3 -8

OpenStudy (cwrw238):

the coordinates of the midpoint of line joining (x1,y1) and (x2,y2) are given by (x1 + x2)/2 and (y1+y2)/2

OpenStudy (cwrw238):

so x-coordinate is (2 + (-6)/2 = -4/2 = -2 can u work out the y-value?

OpenStudy (anonymous):

yes, thanks, i got what u did.

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

I did what u did, my points that i got was (-2,7) would my answer be 7?